# Xuezhang Chen  and Shihong Zhang

### v1

@@ 4 Biharmonic Green function
For $n\geq 2$, we let $X=(x,t)\in\mathbb{R}_{+}^{n+1},Y=(y,s)\in\mathbb{R}_{+}^{n+1}$. We define $\bar{X}=(x,-t)$, and $\xi=F(X),\eta=F(Y)$ with $F$ being the conformal map as in (3.7).
4.1 Noncritical dimension
Recall that the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$ is given by
@@ 4.1 Noncritical dimension
Proof of Theorem 1.3. We denote by $G^{(i,j)}$ and $\bar{G}^{(i,j)}$ the Green functions in +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, respectively.
Fix $X\in\mathbb{R}_{+}^{n+1}$, then for all $Y\in\partial\mathbb{R}_{+}^{n+1}$, direct computations (cf. calculations in the proof of Theorem 1.1) show
@@
The above formulae demonstrate an intimate connection between the Poisson kernel and fundamental function.
By definition of Green functions on both +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, we invoke the conformal invariance of $P_{4}^{g}$ to show
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}$ (4.2)
by setting $u=U\circ FU_{0}$ and $f=\bar{f}\circ FU_{0}^{(n+5)/(n-3)}$. For brevity, we let
@@
$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1}.\end{cases}$
@@
Let $\xi^{\ast}=|\xi|^{-2}\xi$ whenever $\xi\neq 0$, then it is not hard to see that $F(\bar{X})=\xi^{\ast}$. Using
@@
Consequently, by (4.2), (3.8) and (4.3) we obtain
$\bar{G}^{(0,2)}(\xi,\eta)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}G^{(0,2)}(X,Y)$
$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
For $(i,j)=(1,3)$, similarly we have
$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(\bar{X}-Y)$
and
$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
For $(i,j)=(0,1),(2,3)$, the Boggio formula (see, Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]) gives
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\int_{1}^{\frac{|\xi||\xi^{\ast}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-n}\mathrm{d}s$
$=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
Set
$f_{2}(y)=-\mathscr{B}^{3}_{2}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{1}(x-y,t)$
and
$f_{3}(y)=-\mathscr{B}^{3}_{3}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{0}(x-y,t).$
We apply Theorem 1.1 together with (4.1) to solve
$=$ $\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z.$
A similar trick gives
$H^{(0,1)}(X,Y)=-H^{(2,3)}(X,Y).$
Hence, by (3.8) and (4.2) we have
$\bar{H}^{(2,3)}(\xi,\eta)=$ $\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(2,3)}(X,Y)$
$=$ $-\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(0,1)}(X,Y)=-\bar{H}^{(0,1)}(\xi,\eta),$
from which we conclude that
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
Until now, we have already obtained biharmonic Green functions on $\mathbb{R}_{+}^{n+1}$ at the same time.
Observe that
$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2}),\qquad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}},$
then it follows that
@@
Consequently, we combine these facts together to conclude the second assertion. $\Box$
@@ 4.2 Critical dimension
The fundamental solution for $\Delta^{2}$ on $\mathbb{R}_{+}^{4}$ is given by
@@
Proof of Theorem 1.5. Fix $X\in\mathbb{R}^{4}_{+}$, then for all $Y\in\partial\mathbb{R}^{4}_{+}$, it is straightforward to show
$\mathscr{B}^{3}_{0}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{3}(x-y,t);$
$\mathscr{B}^{3}_{1}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{2}(x-y,t);$
$\mathscr{B}^{3}_{2}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{1}(x-y,t);$
$\mathscr{B}^{3}_{3}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{0}(x-y,t).$
We next derive the Green function on $\mathbb{B}^{4}$ associated with the unique smooth solution of
@@
As before, we let
$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
satisfy
$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+}.\end{cases}$
For $(i,j)=(0,2)$, we find
$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y)$
whence,
$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
This together with (3.8), (4.3) gives
$\bar{G}^{(0,2)}(\xi,\eta)=\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|).$
We introduce
$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)+C,\quad C\in\mathbb{R}.$
For $(i,j)=(0,1),(2,3)$, again the Boggio formula gives
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{3}|}\int_{1}^{\frac{|\xi||\xi^{*}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-3}\mathrm{d}s$
$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right).$
Notice that
$=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z+\frac{C}{2},$
then
$H^{(2,3)}(X,Y)=-H^{(0,1)}(X,Y)+\frac{C}{2}.$
Consequently, we have
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left[\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right]+C.$
$\Box$
Eventually, combining the biharmonic Poisson kernel and Green function, we are in a position to present a Green formula for an nonhomogeneous biharmonic boundary value problem on $\mathbb{B}^{n+1}$.
Corollary 4.1 (Green formula for a biharmonic boundary value problem).
@@
As an application of the Green formula, we are able to obtain the following comparison principles for biharmonic boundary value problems on $\mathbb{B}^{n+1}$.
@@
This is a direct consequent of Corollary 4.1, Theorem 1.2 and Theorem 1.3. $\Box$
@@ 5 Classification theorems
As above, let $P_{k}(X)$ and $P_{k}(x)$ denote a polynomial with degree $k$ in the variables $X$ and $x$, where $k\in\mathbb{N}$.
@@
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
@@
The classification theorem (see Chen-Li-Ou [9] and Y. Y. Li [24]) for the integral equation above gives
@@
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{in}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=-\frac{n-3}{2}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}&\mathrm{on}\quad\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}u=\frac{(n^{2}-1)(n-3)}{4}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\quad&\mathrm{on}\quad\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.3)
@@
$M_{1}(x,t)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\mathrm{d}y$
$-\frac{n-3}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}\mathrm{d}y.$
@@
$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
@@
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.6)
@@
$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{2}t^{2}.$ (5.7)
Eventually, we combine (5.4), (5.5) and (5.7) to conclude that $c_{2}^{\prime}=c_{2}$.
(2) If $p_{3}<p_{3}^{\ast}$, then again by Lemma 5.1 we obtain $u(x,0)=0$. In this case, $u$ also solves the same homogeneous biharmonic boundary value problem (5.6), thanks to Lemma 2.3, we conclude that $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$. $\Box$

### v2

@@ 4 Biharmonic Green function
For $n\geq 2$, let $X=(x,t)\in\mathbb{R}_{+}^{n+1},Y=(y,s)\in\mathbb{R}_{+}^{n+1}$. We define $\bar{X}=(x,-t)$, and $\xi=F(X),\eta=F(Y)$ with $F$ being the conformal map as in (3.7).
We first introduce a biharmonic Green function on $\mathbb{B}^{n+1}$ with each of four proper conformal boundary operator pairs. Consider an nonhomogeneous biharmonic BVP with vanishing boundary conditions
$\begin{cases}\Delta^{2}U=\bar{f}\qquad\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n}.\end{cases}$ (4.1)
Definition 4.1 (Biharmonic Green function).
Let $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ and $i,j\in\{0,1,2,3\},i+j\neq 3$. A function $\bar{G}^{(i,j)}:(\xi,\eta)\in\overline{\mathbb{B}^{n+1}}\times\overline{\mathbb{B}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (4.1) is called a Green function, if $\bar{G}^{(i,j)}$ satisfies
We notice that $\bar{G}^{(0,1)}$ coincides with the Boggio formula ( e.g., Boggio [4] or Gazzola-Grunau-Sweers [20, Lemma 2.27]). We next combine this and the biharmonic Poisson kernel to derive explicit formulae for other Green functions.
4.1 Noncritical dimensions
Recall the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$:
@@ 4.1 Noncritical dimensions
For brevity, we let $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{n+1}}$. Notice that
$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2})=|\eta|^{2}|\eta^{\ast}-\xi|^{2}$
can be extended to define for all $\xi,\eta\in\mathbb{B}^{n+1}$.
Theorem 4.1.
For $n\neq 3$, the representation formulae of Green functiona $\bar{G}^{(i,j)}$ to (4.1) are
for all $\xi,\eta\in\mathbb{B}^{n+1},\xi\neq\eta$, where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{4}}$. Moreover, there holds
$0\leq\bar{G}^{(0,1)}(\xi,\eta)\leq\bar{G}^{(0,2)}(\xi,\eta)\leq\bar{G}^{(1,3)}(\xi,\eta)\leq\bar{G}^{(2,3)}(\xi,\eta),\quad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}},\xi\neq\eta,$
Proof of Theorems 1.2 and 4.1. Fix $X\in\mathbb{R}_{+}^{n+1}$. Then for all $Y\in\partial\mathbb{R}_{+}^{n+1}$, direct computations (cf. calculations in the proof of Theorem 1.1) show
@@
Here each $\mathscr{B}^{3}_{k},k\in\{1,2,3\}$ acts on $\Gamma$ with respect to the variable $Y$.
By definition of Green functions on +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, we invoke the conformal invariance of $P_{4}^{g}$ to show
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}.$ (4.3)
For brevity, we let
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
@@
For $(i,j)=(1,3)$, similarly we have
$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(\bar{X}-Y).$
For $(i,j)=(0,1),(2,3)$, the Boggio formula (see, Boggio [4] or Gazzola-Grunau-Sweers [20, Lemma 2.27]) gives
$G^{(0,1)}(X,Y)=$ $\frac{1}{4|\mathbb{S}^{n}|}|X-Y|^{3-n}\int_{1}^{\frac{|\bar{X}-Y|}{|X-Y|}}(s^{2}-1)s^{-n}\mathrm{d}s$
$=$ $\Gamma(X-Y)+\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right].$
Set
$f_{2}(y)=-\mathscr{B}^{3}_{2}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{1}(x-y,t)$
and
$f_{3}(y)=-\mathscr{B}^{3}_{3}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{0}(x-y,t).$
Thanks to Theorem 1.1 and (4.2), a clever observation is
$=$ $\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z$
$=$ $\int_{\mathbb{R}^{n}}P^{3}_{0}(x-z,t)\mathscr{B}^{3}_{0}\big{(}\Gamma(y-z,s)\big{)}\mathrm{d}z+\int_{\mathbb{R}^{n}}P^{3}_{1}(x-z,t)\mathscr{B}^{3}_{1}\big{(}\Gamma(y-z,s)\big{)}\mathrm{d}z$
$=$ $-H^{(0,1)}(X,Y).$
Then it follows
$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)-\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right].$
Observe that
$|\bar{X}-Y|^{2}=|X-Y|^{2}+4st,\qquad\forall~{}X,Y\in\overline{\mathbb{R}_{+}^{n+1}}.$
Then we have
$G^{(2,3)}(X,Y)\geq\Gamma(X-Y)+\Gamma(|\bar{X}-Y|)=$ $G^{(1,3)}(X,Y)\geq G^{(0,2)}(X,Y)\geq 0,$
$[G^{(0,2)}-G^{(0,1)}](X,Y)=$ $\frac{n-3}{2}\Gamma(|\bar{X}-Y|)\left(1-\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}\right)\geq 0.$
This finish the proof of Theorem 1.2.
Let $\xi^{\ast}=|\xi|^{-2}\xi$ whenever $\xi\neq 0$. Noitice that $F(\bar{X})=\xi^{\ast}$. Using
@@
Therefore, using (4.3), (3.8) and (4.4) the explicit representation formulae $\bar{G}^{(i,j)}$ for $\mathbb{B}^{n+1}$ follow from those of $G^{(i,j)}$ for $\mathbb{R}_{+}^{n+1}$.
Notice that
$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2}),\qquad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}}.$
Then
@@
Hence the second assertion follows. This completes the proof of Theorem 4.1. $\Box$
@@ 4.2 Critical dimension
The purpose is to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
Recall that the fundamental solution for $\Delta^{2}$ on $\mathbb{R}_{+}^{4}$ is
@@
Theorem 4.2.
The representation formula of Green function $\bar{G}^{(i,j)}$ to (4.1) is
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right)+C;$
$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+C;$
for $\xi,\eta\in\overline{\mathbb{B}^{4}},\xi\neq\eta$ and $C\in\mathbb{R}$.
Proof of Theorems 1.3 and 4.2. By definition of biharmonic Green function, we have
$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
satisfy
$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{~{}~{}in~{}~{}}\qquad~{}~{}\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{4}_{+}.\end{cases}$
For $(i,j)=(0,2)$, we find
$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y)$
whence,
$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
By (1.7) and symmetry we find
$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(|\bar{X}-Y|)+C,\quad C\in\mathbb{R}.$
For $(i,j)=(0,1),(2,3)$, again the Boggio formula gives
$G^{(0,1)}(X,Y)=$ $\frac{1}{4|\mathbb{S}^{3}|}\int_{1}^{\frac{|\bar{X}-Y|}{|X-Y|}}(s^{2}-1)s^{-3}\mathrm{d}s$
$=$ $\Gamma(X-Y)-\Gamma(|\bar{X}-Y|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right).$
Notice that
$=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z+\frac{C}{2}.$
Then
$H^{(2,3)}(X,Y)=-H^{(0,1)}(X,Y)+\frac{C}{2}.$
Consequently, we have
$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(|\bar{X}-Y|)-\frac{1}{8|\mathbb{S}^{3}|}\left[\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right]+C.$
This finishes the proof of Theorem 1.3.
Next, we derive the Green function on $\mathbb{B}^{4}$ associated with the unique smooth solution of
@@
This completes the proof of Theorem 4.2. $\Box$
Combining the biharmonic Poisson kernel and Green function, we are in a position to present a Green formula for an nonhomogeneous biharmonic BVP on $\mathbb{B}^{n+1}$.
Corollary 4.1 (Biharmonic Poisson and Green integral).
@@
As an application of the Green formula, we are able to obtain the following comparison principles for biharmonic BVPs on $\mathbb{B}^{n+1}$.
@@
This is a direct consequence of Corollary 4.1, Theorems 3.1 and 4.1. $\Box$
@@ 5 Classification theorems
Our approach builds on the intrinsic third-order GJMS operator $P_{3}=\frac{1}{2}\mathscr{B}_{3}^{3}$ in our setting, and the basics of the geometric bubble and biharmonic Poisson integral.
An important property of the geometric bubble $U_{x_{0},\varepsilon}$ is that for $n\geq 4$ and $i,j\in\{1,2,3\},i<j$, $U_{x_{0},\varepsilon}$ satisfies the BVP (1.5) with critical growth that $p_{i}=p_{i}^{\ast},p_{j}=p_{j}^{\ast}$.
There are two ways to see this. One is through a straightforward computation. The other is in a geometric way of exhausting all possible geometric bubbles: A conformal transformation on $\overline{\mathbb{B}^{n+1}}$ can be expressed as
$\zeta:=\psi_{a}(\xi)=\frac{\xi-a-|\xi|^{2}a+2(a\cdot\xi)a-|a|^{2}\xi}{1-2a\cdot\xi+|a|^{2}|\xi|^{2}},\quad\forall~{}\xi\in\mathbb{B}^{n+1}$
for some $a\in\mathbb{B}^{n+1}$ (see e.g., [28, Chapter 1]), and there holds
$\psi_{a}^{\ast}(|\mathrm{d}\zeta|^{2})=\left(\frac{1-|a|^{2}}{|\xi|^{2}|a|^{2}-2a\cdot\xi+1}\right)^{2}|\mathrm{d}\xi|^{2}.$
Indeed, one can verify that
$F^{\ast}(\psi_{a}^{\ast}(|\mathrm{d}\zeta|^{2}))=U_{x_{0},\varepsilon}^{\frac{4}{n-3}}(X)|\mathrm{d}X|^{2}$
through the relation
$a=F(x_{0},\varepsilon)\qquad\mathrm{by~{}~{}regarding~{}~{}}(x_{0},\varepsilon)\in\mathbb{R}_{+}^{n+1}.$
Equivalently, the BVP (1.5) with critical growth can interpreted as the pullback of the metric $(\mathbb{B}^{n+1},\psi_{a}^{\ast}(|\mathrm{d}\zeta|^{2}))$ via the map $F$ has constant curvatures $(T_{i}^{3})_{\mathbb{S}^{n}},i\in\{1,2,3\}$.
As before, let $P_{k}(x)$ denote a polynomial with degree $\leq k$ for $k\in\mathbb{N}$.
@@
$\begin{cases}\displaystyle~{}\Delta^{2}w=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}~{}\partial_{t}w=0&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
@@
The classification theorem (e.g., Y. Y. Li [30] and Chen-Li-Ou [12]) for the integral equation above gives
@@
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}\partial_{t}u=-\frac{n-3}{2}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\Big{)}^{\frac{(n-3)p_{1}}{2}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}u=\frac{(n^{2}-1)(n-3)}{4}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\Big{)}^{\frac{n+3}{2}}\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.3)
@@
$M_{1}(x,t)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\Big{)}^{\frac{n+3}{2}}\mathrm{d}y$
$-\frac{n-3}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\Big{)}^{\frac{(n-3)p_{1}}{2}}\mathrm{d}y.$
@@
$w(x,t)=u(x,t)-U_{x_{0},\varepsilon}(x,t),$
@@
$\begin{cases}\displaystyle~{}\Delta^{2}w=0\qquad\qquad\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}~{}\partial_{t}w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.6)
@@
$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{2}t^{2}.$ (5.7)
Therefore, we combine (5.4), (5.5) and (5.7) to conclude that $M_{1}=U_{x_{0},\varepsilon}$ and $c_{2}^{\prime}=c_{2}$.
(2) If $p_{3}<p_{3}^{\ast}$, then again by Lemma 5.1 we obtain $u(x,0)=0$. In this case, $u$ also solves the same homogeneous biharmonic BVP (5.6), thanks to Lemma 2.3, we conclude that $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$. $\Box$
@@
From the viewpoint of the Lopatinski-Shapiro or complement condition, the conformal boundary operator pair $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$ is exceptional. Nevertheless, some different technique needs to be developed for treating this type of $T$-curvature equations.